OpenStudy (anonymous):

22/7 divided by 11/14

OpenStudy (anonymous):

(22/7)/(11/14)=2/(1/2)=4

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

BA?

OpenStudy (anonymous):

4